OpenStudy (anonymous):

evaluate the iterated integral ∫0 to 4 ∫ (−sqrt(16−x^2) to (√16−x2) e^(−x^2−y^2) dy dx . Also, sketch the planar region R that is determined by the limits of the integration.

OpenStudy (anonymous):

evaluate the iterated integral \[\int\limits_{0}^{4}\int\limits_{-\sqrt{16-x^2}}^{\sqrt{16-x^2}} e ^{-x^2-y^2} dy dx \] . Also, sketch the planar region R that is determined by the limits of the integration.